# Comparison of modularity-based approaches for nodes clustering in binary hypergraphs

### 25 Jan 2024 v1

@@ Appendix B All-or-nothing modularity revisited
where $\hat{\beta}_{s}=\log\hat{\omega}_{s1}-\log\hat{\omega}_{s0}$ and $\hat{\gamma}_{s}=\hat{\beta}_{s}^{-1}(\hat{\omega}_{s1}-\hat{\omega}_{s0})$, the cut terms $\mathrm{cut}_{s}(\mathcal{C})$ count the (weighted) number of hyperedges of size $s$ that contain nodes in two or more distinct clusters; namely, $\mathrm{cut}_{s}(\mathcal{C})=|\mathcal{E}_{s}|-e_{H}(C_{(s)})$. The first line in Equation 11 is the expression of $Q^{\textrm{aon}}$ given in Chodrow et al. (2021) while the following lines correspond to our rewriting, showing the similarities with the other previously defined modularities. While in general we may expect that $\hat{\omega}_{s1}>\hat{\omega}_{s0}$ so that both $\hat{\beta}_{s},\hat{\gamma}_{s}>0$, we then recover in this expression a sum of difference terms between a count of specific hyperedges (those entirely included in a cluster) and a correcting volume term.

### 04 Mar 2024 v2

@@ Appendix B All-or-nothing modularity revisited
where $\hat{\beta}_{s}=\log\hat{\omega}_{s1}-\log\hat{\omega}_{s0}$ and $\hat{\gamma}_{s}=\hat{\beta}_{s}^{-1}(\hat{\omega}_{s1}-\hat{\omega}_{s0})$, the cut terms $\mathrm{cut}_{s}(\mathcal{C})$ count the (weighted) number of hyperedges of size $s$ that contain nodes in two or more distinct clusters; namely, $\mathrm{cut}_{s}(\mathcal{C})=|\mathcal{E}_{s}|-e_{H}(C_{(s)})$. The first line in Equation (11) is the expression of $Q^{\textrm{aon}}$ given in Chodrow et al. (2021) while the following lines correspond to our rewriting, showing the similarities with the other previously defined modularities. While in general we may expect that $\hat{\omega}_{s1}>\hat{\omega}_{s0}$ so that both $\hat{\beta}_{s},\hat{\gamma}_{s}>0$, we then recover in this expression a sum of difference terms between a count of specific hyperedges (those entirely included in a cluster) and a correcting volume term.
@@ Appendix C Linking parameters in HSBM and DCHSBM-like
Appendix D Values for ground truth modularity
Table 5 shows mean and standard deviations values of the modularities at the ground truth clustering for each method (each one focusing on a specific modularity definition, as summarized in Table 1) and each scenario. We observe that the IRMM algorithm that maximizes $Q^{\textrm{w-clique}}$ has a ground truth modularity close to 0.
Table 5: Mean value (and standard deviation) of ground truth modularities $Q^{\textrm{true}}=Q(H,\mathcal{C}^{\textrm{true}})$ for the 25 hypergraphs generated under each method and scenario.
scenario / method AON-HMLL CNM IRMM LSR
ScenA-HSBM
A1: $n=50$1440.28(253.46) 0.31(0.04) -0.01(0.02) 0.23(0.04)
A2: $n=100$3281.04(321.45) 0.34(0.02) -0.01(0.02) 0.27(0.02)
A3: $n=150$5089.41(349.39) 0.35(0.02) 0(0.01) 0.27(0.02)
A4: $n=200$7007.19(431.96) 0.36(0.02) 0(0.01) 0.28(0.02)
A5: $n=500$19151.93(657.66) 0.36(0.01) 0(0.01) 0.28(0.01)
ScenA-DCHSBM
A1: $n=50$1507.05(140.8) 0.38(0.03) -0.02(0.02) 0.3(0.03)
A2: $n=100$3214.43(259.46) 0.37(0.02) 0(0.02) 0.29(0.02)
A3: $n=150$5055.25(307.23) 0.37(0.02) -0.01(0.01) 0.29(0.02)
A4: $n=200$6819.27(297.63) 0.37(0.02) 0(0.01) 0.29(0.02)
A5: $n=500$18619.63(568.68) 0.36(0.01) 0(0.01) 0.29(0.01)
A6: $n=1000$39925.19(782.67) 0.36(0.01) 0(0.01) 0.29(0.01)
ScenA-hABCD
A1: $n=50$177.69(39.81) 0.27(0.1) -0.02(0.07) 0.24(0.09)
A2: $n=100$439.39(58.81) 0.34(0.07) -0.01(0.04) 0.3(0.06)
A3: $n=150$727.05(63.2) 0.36(0.04) -0.01(0.03) 0.31(0.03)
A4: $n=200$1040.2(69.62) 0.38(0.04) -0.01(0.03) 0.33(0.04)
A5: $n=500$3453.7(312.77) 0.42(0.01) 0(0.02) 0.36(0.01)
A6: $n=1000$8578.86(633.97) 0.44(0.01) 0(0.01) 0.37(0.01)
ScenB-DCHSBM
B1: $n=50$4476.24(286.51) 0.38(0.02) -0.01(0.01) 0.3(0.02)
B2: $n=100$9895.18(438.98) 0.37(0.01) -0.01(0.01) 0.29(0.01)
B3: $n=150$15436.71(407.27) 0.37(0.01) 0(0.01) 0.29(0.01)
B4: $n=200$21146.49(497.66) 0.36(0.01) 0(0.01) 0.29(0.01)
B5: $n=500$57396.92(1005.76) 0.37(0) 0(0) 0.29(0)
B6: $n=1000$122253.16(1071.73) 0.36(0) 0(0) 0.29(0)
ScenC-HSBM
C1: $n=50$1692.96(341.87) 0.22(0.04) -0.01(0.02) 0.16(0.04)
C2: $n=100$3818.35(526.14) 0.23(0.03) -0.01(0.01) 0.16(0.03)
C3: $n=150$6006.44(720.38) 0.24(0.03) 0(0.01) 0.17(0.03)
C4: $n=200$8411.32(678.46) 0.24(0.02) 0(0.01) 0.17(0.02)
C5: $n=500$23301.14(1133.75) 0.24(0.02) 0(0.01) 0.17(0.02)
ScenD-DCHSBM
D1: $n=50$3268.27(172.98) 0.47(0.03) -0.01(0.02) 0.28(0.02)
D2: $n=100$7377.67(213.42) 0.45(0.02) 0(0.01) 0.27(0.01)
D3: $n=150$11692.36(325.51) 0.45(0.02) -0.01(0.01) 0.27(0.01)
D4: $n=200$15946(274.87) 0.45(0.01) 0(0.01) 0.27(0.01)
D5: $n=500$43675.15(431.82) 0.45(0.01) 0(0.01) 0.27(0.01)
D6: $n=1000$92887.89(1079.8) 0.45(0.01) 0(0) 0.27(0.01)
ScenE-DCHSBM
E1: $n=50$1496.2(102.41) 0.41(0.03) -0.01(0.02) 0.33(0.02)
E2: $n=100$3210.91(190.09) 0.41(0.02) 0(0.01) 0.33(0.02)
E3: $n=150$5117.84(249) 0.41(0.02) -0.01(0.01) 0.32(0.02)
E4: $n=200$6902.04(311.24) 0.4(0.02) -0.01(0.01) 0.32(0.01)
E5: $n=500$19069.69(632.72) 0.4(0.01) 0(0.01) 0.32(0.01)
E6: $n=1000$40020.25(699.77) 0.4(0) 0(0) 0.32(0)
ScenF-DCHSBM
F1: $n=50$1401.74(122.37) 0.33(0.03) -0.02(0.02) 0.26(0.02)
F2: $n=100$3045.16(240.8) 0.32(0.02) 0(0.02) 0.25(0.02)
F3: $n=150$4990.03(275.21) 0.32(0.02) 0(0.01) 0.25(0.01)
F4: $n=200$6623.92(208.33) 0.33(0.02) 0(0.01) 0.26(0.02)
F5: $n=500$18654.39(359.67) 0.32(0.01) 0(0.01) 0.25(0.01)
F6: $n=1000$39341.82(941.03) 0.32(0.01) 0(0) 0.25(0.01)
ScenZ-hABCD
Z1: $n=100$529.92(164.11) 0.28(0.07) -0.02(0.04) 0.26(0.08)
Z2: $n=150$782.78(209.52) 0.33(0.05) -0.02(0.04) 0.31(0.05)
Z3: $n=200$961.22(128.09) 0.35(0.04) -0.01(0.02) 0.34(0.03)
Z4: $n=500$2457.49(132.12) 0.39(0.02) 0(0.01) 0.41(0.02)
Z5: $n=1000$5027.19(177.84) 0.41(0.01) 0(0.01) 0.43(0.01)